Write an equation in slope-intercept form of each line described.
Slope = 2; passes through (2, -1)

Answer by mananth(16946)

Slope = 2; passes through (2, -1)
equation of line
(y-y1)=m(x-x1)
(x1,y1)(2,-1)
m=-2
(y-(-1))=-2(x-2)
y+1=-2x+4
y=-2x+3